A) If a water balloon is thrown straight up at 12 m/s from the top of a 13 m building, how long will it take to hit the ground?

B) If a water balloon is thrown straight down at 14 m/s from the top of a 31 m building, how long will it take to hit the ground?

C) If a water balloon is thrown upwards at a 17 degree angle at 14 m/s from the top of a 31 m building, how far from the base of the building will it hit (hint, first find how long will it take to hit the ground).

D) If a water balloon is thrown downwards at a 18 degree angle (measured from the horizontal) at 7 m/s from the top of a 20 m building, how far from the base of the building will it hit (hint, first find how long will it take to hit the ground).
Answer #1

a)

usimg 2nd equation of motion

0 = h + u t - 0.5 gt^2

0 = 13 + 12* t - 4.9 t^2

solve for t

t = 3.262 s

=======

b)

similarly

31 = 14 t + 4.9 t^2

solve for t

t = 1.464 s

=====

c)

motion along vertical

0 = 31 + 14* sin 17 t - 4.9* t^2

solve for t

t = 2.967 s

motion along horizontal

R = 14 cos 17 * t =39.72 m

========

d)

motion along vertical

0 = 20 + 7 sin 18 t - 4.9* t^2

solve for t

t = 2.253 s

motion along horizontal

R = 7 cos 18 * t = 15 m

======

Comment before rate in case any doubt, will reply for sure.. goodluck